OpenStudy (anonymous):

Solve sin(11x)=1 on the interval [0, 2pi/11]

OpenStudy (anonymous):

11x=pi/2 x=pi/22

OpenStudy (anonymous):

=> 11x=npi/2 wer n is 4m+1 m=0,1,2.....

OpenStudy (anonymous):

pi/22 is the solution

OpenStudy (anonymous):

that's what i got sweet!